# Using Dynamic Safety Margins as Control Barrier Functions

### v3

@@ 1 Introduction
Given a state constraint set with a smooth boundary, it is usually straightforward to find a differentiable function of the state vector that is positive for constraint-admissible states, and negative otherwise. We call such functions candidate CBFs. They are commonly used in practice to design CBF-based controllers and, with tuning, can achieve great performance (e.g., [2]). However, candidate CBFs lack many safety guarantees that valid CBFs provide. In particular, the CBF-based optimization program governing the control law may become infeasible when using candidate CBFs, which is unacceptable for safety-critical systems. Ultimately, CBFs are certificates of control invariance. Therefore, if the original state constraint set is not control invariant, it is impossible to find a valid CBF associated to it. In these cases, we must resort to finding CBFs for a, hopefully large, control invariant subset of the state constraint set. In this work, we use tools from the reference governor literature [4] to systematically design valid CBFs given arbitrary state and input constraints.
Figure 1: Comparison of CBF-based control (top), which filters a nominal input signal $\boldsymbol{\mathbf{u}}_{n}$ to obtain a safe input $\boldsymbol{\mathbf{u}}^{*}$, with a specific reference governor approach, the command governor (CG). The CG filters the desired reference signal $\boldsymbol{\mathbf{r}}$ to obtain a safe virtual reference $\boldsymbol{\mathbf{v}}^{*}$ for the prestabilizing controller.
@@ 1.1 Related Works
Most relevant to the present manuscript, we review some works that study the design of CBFs and feasibility of the CBF-based optimization program. In [14], sum of squares (SOS) techniques are used to synthesize CBFs for semi-algebraic sets and polynomial systems using an alternating descent algorithm. In [15], the authors use SOS techniques to synthesize polynomial control Lyapunov functions, and then present an extension to construct CBFs. In [16], the authors introduce a variable modifying the decay rate of the CBF to guarantee feasibility of the CBF-based program when the state is in the interior of the safe set even under input constraints. However, their approach provides no feasibility guarantee when the state is on the boundary of the safe set. In [17], the authors use control Lyapunov functions (CLFs) to design a continuum of valid CBFs and use tools from hybrid systems theory to study the resulting safety properties. In [18], the authors design CBFs for Euler–Lagrange systems. While the results are promising and the class of systems is relevant, the approach is limited to box constraints. In [19], the authors learn CBFs from human demonstrations but they provide no guarantees that the learned CBFs are valid. In [20], given a valid CBF, the authors use learning to construct a CBF for a larger control invariant set using a combination of safe and unsafe trajectories of the system. Nonetheless, finding the initial, valid CBF remains an open question. In [21], the authors use a backup control policy to enlarge a small (but easy to find) control invariant set. The enlarged control invariant set is then used to derive a backup CBF. These backup CBFs rarely have a closed form, which makes them difficult to implement. In [22], the authors use maximal output admissible sets (MOASs) to design discrete-time CBFs for arbitrary state and input constraints. While the approach works well for linear systems, finding the MOAS for nonlinear systems remains an open question.
@@ 1.2 Contributions
  • We show that dynamic safety margins (DSMs) are (vector-valued) CBFs for an augmented system consisting of the concatenation of the state vector and the reference of the prestabilizing controller used to construct the DSM.
  • We show that the DSM-CBFs support constraints with arbitrary relative degree.
  • We formulate the DSM-CBF optimization-based policy and show that it is safe and persistently feasible. Moreover, the optimization problem is a quadratic program (QP) when the input constraints are polyhedral.
  • We provide conditions that guarantee the DSM-CBF optimization-based policy is locally Lipschitz continuous. This guarantees existence and uniqueness of the closed-loop trajectory.
  • We extend the Lyapunov-based DSMs presented in [23] to inadmissible references. This extension is needed to use them as DSM-CBFs. Moreover, we show that any class $\mathscr{K}_{\infty}$ function can be used with Lyapunov-based DSM-CBFs.
  • We present a comparison of our proposed approach with the explicit reference governor (ERG) [23], candidate CBFs and backup CBFs [21] in two nonlinear systems. The numerical simulations show that the DSM-CBFs we propose outperform the ERG, while boasting strict safety guarantees that are otherwise lost when candidate CBF-based control becomes infeasible. The examples also show that, while backup CBFs can achieve better performance than DSM-CBFs, their implementation is challenging in the presence of multiple constraints or loss of control authority.
@@ 1.3 Notation
Throughout the paper, only vector quantities are presented in bold font: $\boldsymbol{\mathbf{x}}\in\mathbb{R}^{n}$. The notation $[\boldsymbol{\mathbf{x}};~{}\boldsymbol{\mathbf{y}}]$ denotes the vertical stacking of two vectors $\boldsymbol{\mathbf{x}}$ and $\boldsymbol{\mathbf{y}}$ into a column vector. A continuous function $\alpha:[0,\infty)\to[0,\infty)$ is said to belong to class $\mathscr{K}_{\infty}$ if it is strictly increasing, $\alpha(0)=0$, and $\lim_{r\to\infty}\alpha(r)=\infty$. The symbol $\subset$ denotes set inclusion, but not necessarily strict inclusion. Given a set $\mathcal{S}$, we denote by $\operatorname{Int}\left(\mathcal{S}\right)$ the interior of $\mathcal{S}$ and by $\partial\mathcal{S}$ the boundary of $\mathcal{S}$ (i.e. $\partial\mathcal{S}=\text{closure}(\mathcal{S})\setminus\operatorname{Int}\left(\mathcal{S}\right)$). Given a subset $\mathcal{C}\subset\mathbb{R}^{n}$ of the state space $\mathbb{R}^{n}$, we generally denote by $\tilde{\mathcal{C}}\subset\mathbb{R}^{n}\times\mathbb{R}^{l}$ its counterpart in the augmented state-reference space. Let $h:\mathcal{D}\to\mathbb{R}$ be continuously differentiable. We say $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ is a regular point of $h$ if $\partial h(\boldsymbol{\mathbf{x}})/\partial\boldsymbol{\mathbf{x}}\neq 0$. We say $c\in\mathbb{R}$ is a regular value of $h$ if every $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ such that $h(\boldsymbol{\mathbf{x}})=c$ is a regular point.
@@ 2.1 Control Invariance
Note that this concept is sometimes called weak control invariance [26] because it doesn’t require existence and uniqueness of the trajectory. Nonetheless, it will simplify the presentation of the upcoming results. Assumptions for the existence and uniqueness of solutions will be presented separately. Central to the upcoming discussion is Nagumo’s theorem characterizing control invariance in terms of Bouligand’s tangent cone.
@@
A compact set $\mathcal{C}\subset\mathbb{R}^{n}$ is control invariant if and only if
@@
where $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is the tangent cone to $\mathcal{C}$ in $\boldsymbol{\mathbf{x}}$ [26, Definition 3.1].
While this characterization was groundbreaking, working with the tangent cone $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is cumbersome. Fortunately, when the compact set $\mathcal{C}$ has a smooth boundary (i.e., can be represented as the superlevel set of a continuously differentiable function), the tangent cone on the boundary is reduced to the tangent half-space shifted to the origin and we have the following characterization.
Theorem 2 ([28, Proposition 4.3.7])
Let $h:\mathcal{D}\to\mathbb{R}$ be a continuously differentiable function with regular value 00 and suppose $\mathcal{C}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}:h(\boldsymbol{\mathbf{x}})\geq 0\}$ is compact. The set $\mathcal{C}$ is control invariant if and only if
@@
Note that [28, Proposition 4.3.7] is more general than Theorem 2: it allows the set $\mathcal{C}$ to be defined using multiple continuously differentiable functions. We use this fact in upcoming results.
@@ 2.2 Control Barrier Functions
Control barrier functions are real-valued functions that certify the control invariance of their superlevel set.
@@
A continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ is a control barrier function (CBF) if there exists a class $\mathscr{K}_{\infty}$ function $\alpha:[0,\infty)\to[0,\infty)$ such that
@@
Assume $\mathcal{U}\!\subset\!\mathbb{R}^{m}$ is compact. Let $\mathcal{C}\!\subset\!\mathbb{R}^{n}$, defined as in (5), be compact and let 00 be a regular value of $h$. The set $\mathcal{C}$ is control invariant if and only if $h$ is a CBF.
@@
is continuous on $\mathcal{D}$. Moreover, it can be shown that the set-valued map $C$ is both uppper- and lower-hemicontinuous. Thus, it follows by Berge’s maximum theorem [29], that the value function $\hat{\alpha}$ is continuous on $[0,\infty)$. By construction, $\hat{\alpha}$ is non-decreasing. By Theorem 2, for any $\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$ such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq 0$. Thus, $\forall\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}=\mathcal{C}(0)$, $\xi(\boldsymbol{\mathbf{x}})\geq 0$. As a result,
@@
There always exists a class $\mathscr{K}_{\infty}$ function $\alpha:[0,\infty)\to[0,\infty)$ that upper-bounds $\hat{\alpha}$. Finally, let $\boldsymbol{\mathbf{x}}\in\mathcal{C}$ be given and note
@@
On the other hand, assume $h$ is a CBF, let $\alpha\in\mathscr{K}_{\infty}$ be such that (4) holds, and let $\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}$ be given. By compactness of $\mathcal{U}$, there exists $\boldsymbol{\mathbf{u}}\in\mathcal{U}$ such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq-\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}=0$, where the last equality holds because $h(\boldsymbol{\mathbf{x}})=0$. Since $\mathcal{C}$ is compact and 00 is a regular value of $h$, we can apply Theorem 2 to conclude $\mathcal{C}$ is control invariant.
@@
In this work we drop the compactness requirement on $\mathcal{U}$ because only the sufficiency statement in Lemma 1 is relevant and we will design CBFs for which (9) holds.
@@
Consider $n_{c}\in\mathbb{N}$ continuously differentiable functions $h_{i}:\mathcal{D}_{i}\to\mathbb{R}$ and let $\mathcal{C}_{i}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}_{i}\mid h_{i}(\boldsymbol{\mathbf{x}})\geq 0\}\subset\mathbb{R}^{n}$, with $\mathcal{C}=\cap_{i=1}^{n_{c}}\mathcal{C}_{i}$. The functions $h_{i}$ are said to have the control-sharing property if there exists a class $\mathscr{K}_{\infty}$ function $\alpha$ satisfying $\forall\boldsymbol{\mathbf{x}}\in\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$, such that
@@
In this paper, we call a smooth function $h:\mathcal{D}\to\mathbb{R}^{n_{c}}$ whose channels $h_{i}$ have the control-sharing property a (vector-valued) control barrier function.
The functions $h_{i}$ in Definition 2.4 are deliberately not labeled as CBFs individually (as opposed to [9, Definition 2]) because such a requirement is unnecessarily limiting. In fact, if 00 is a regular value of each $h_{i}$, Lemma 1 also holds for the set $\mathcal{C}$ defined as the intersection of the $\mathcal{C}_{i}$ even if each $\mathcal{C}_{i}$ is individually not compact (see [28, Proposition 4.3.7] for more details.) The optimization-based policy $\beta$ can also be used with (vector-valued) CBFs with the following admissible input set
@@
As noted in the introduction, finding a CBF for arbitrary constraint sets $\mathcal{X}\subset\mathbb{R}^{n}$ is a challenging task. In fact, Lemma 1 states that, when $\mathcal{U}$ is compact, a CBF $h:\mathcal{D}\to\mathbb{R}$ satisfying $\mathcal{X}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}$ with a well-defined policy $\beta$ exists only if $\mathcal{X}$ is control invariant, which is generally not the case. In spite of this, many practitioners insist on imposing $\mathcal{C}=\mathcal{X}$ to construct a “candidate” CBF (i.e. a continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ that satisfies $\mathcal{X}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}$) and then tuning a class $\mathscr{K}_{\infty}$ function $\alpha$ on a case-by-case basis. Although this approach can yield good performance, it suffers from the possibility that $\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})=\emptyset$ for some $\boldsymbol{\mathbf{x}}\in\mathcal{X}$, which makes the controller $\beta$ ill-defined.

### v4

@@ 1 Introduction
Given a state constraint set, it is usually straightforward to find a function of the state vector that is positive for constraint-admissible states, and negative otherwise. We call such functions candidate CBFs. They are commonly used in practice to design CBF-based controllers and, with tuning, can achieve great performance (e.g., [2]). However, candidate CBFs lack many safety guarantees that valid CBFs provide. In particular, the CBF-based optimization program governing the control law may become infeasible when using candidate CBFs, which is unacceptable for safety-critical systems. Ultimately, CBFs are certificates of control invariance. Therefore, if the original state constraint set is not control invariant, it is impossible to find a valid CBF associated to it. In these cases, we must resort to finding CBFs for a, hopefully large, control invariant subset of the state constraint set. In this work, we use tools from the reference governor literature [4] to systematically design valid CBFs given arbitrary state and input constraints.
Figure 1: Comparison of CBF-based control (top) with RG-based control (bottom). The CBF approach filters a nominal input signal $\boldsymbol{\mathbf{u}}_{n}$ to obtain a safe input $\boldsymbol{\mathbf{u}}$. The RG approach a nominal reference signal $\boldsymbol{\mathbf{r}}_{n}$ to obtain a safe virtual reference $\boldsymbol{\mathbf{r}}$ for the nominal controller.
@@ 1.1 Related Works
Most relevant to the present manuscript, we review some works that study the design of CBFs and feasibility of the CBF-based optimization program. In [14], sum of squares (SOS) techniques are used to synthesize CBFs for semi-algebraic sets and polynomial systems using an alternating descent algorithm. In [15], the authors use SOS techniques to synthesize polynomial control Lyapunov functions, and then present an extension to construct CBFs. In [16], the authors introduce a variable modifying the decay rate of the CBF to guarantee feasibility of the CBF-based program when the state is in the interior of the safe set even under input constraints. However, their approach provides no feasibility guarantee when the state is on the boundary of the safe set. In [17], the authors propose robust control barrier-value functions, which unify the Hamilton–Jacobi reachability and CBF methods, leading to viability kernel-sized invariant sets and CBF-like performances. This method, however, suffers from the curse of dimensionality. In [18], the authors use control Lyapunov functions (CLFs) to design a continuum of valid CBFs and use tools from hybrid systems theory to study the resulting safety properties. In [19], the authors design CBFs for Euler–Lagrange systems. While the results are promising and the class of systems is relevant, the approach is limited to box constraints. In [20], the authors learn CBFs from human demonstrations but they provide no guarantees that the learned CBFs are valid. In [21], given a valid CBF, the authors use learning to construct a CBF for a larger control invariant set using a combination of safe and unsafe trajectories of the system. Nonetheless, finding the initial, valid CBF remains an open question. In [22], the authors use a backup control policy to enlarge a small (but easy to find) control invariant set. The enlarged control invariant set is then used to derive a backup CBF. These backup CBFs rarely have a closed form, which makes them difficult to implement. In [23], the authors use maximal output admissible sets (MOASs) to design discrete-time CBFs for arbitrary state and input constraints. While the approach works well for linear systems, finding the MOAS for nonlinear systems remains an open question.
@@ 1.2 Contributions
  • We show that dynamic safety margins (DSMs) are (vector-valued) CBFs for an augmented system consisting of the concatenation of the state vector and the reference of the prestabilizing controller used to construct the DSM. Additionally, we show that the relative degree of the constraints is irrelevant for DSMs by construction.
  • We formulate the CBF safety filter in the augmented state-reference space and show that it is persistently feasible. Moreover, the optimization problem is a quadratic program (QP) when the input constraints are polyhedral.
  • We extend the Lyapunov-based DSMs presented in [24] to inadmissible references. This extension is needed to use them as (vector-valued) CBFs. Moreover, we show that any class $\mathscr{K}$ function satisfies the CBF condition of Lyapunov-based DSMs.
  • We present a comparison of our proposed approach with the explicit reference governor (ERG) [24], candidate CBFs and backup CBFs [22] in two nonlinear systems. The numerical simulations show that the proposed DSM-CBFs outperform the ERG, while boasting strict safety guarantees that are otherwise lost when candidate CBF-based control becomes infeasible. The examples also show that, while backup CBFs can achieve better performance than Lyapunov-based DSM-CBFs, their implementation is more computationally expensive due to its reliance on trajectory predictions.
@@ 1.3 Notation
Throughout the paper, only vector quantities are presented in bold font: $\boldsymbol{\mathbf{x}}\in\mathbb{R}^{n}$. The notation $\boldsymbol{\mathbf{x}};~\boldsymbol{\mathbf{y}}]$ denotes the vertical stacking of two vectors $\boldsymbol{\mathbf{x}}$ and $\boldsymbol{\mathbf{y}}$ into a column vector. A continuous function $\alpha:[0,\infty)\to[0,\infty)$ is said to belong to class $\mathscr{K}$ if it is strictly increasing and $\alpha(0)=0$. The symbol $\subset$ denotes set inclusion, but not necessarily strict inclusion. Given a set $\mathcal{S}$, we denote by $\operatorname{Int}\left(\mathcal{S}\right)$ the interior of $\mathcal{S}$ and by $\partial\mathcal{S}$ the boundary of $\mathcal{S}$ (i.e. $\partial\mathcal{S}=\text{closure}(\mathcal{S})\setminus\operatorname{Int}\left(\mathcal{S}\right)$). Given a subset $\mathcal{C}\subset\mathbb{R}^{n}$ of the state space $\mathbb{R}^{n}$, we generally denote by $\tilde{\mathcal{C}}\subset\mathbb{R}^{n}\times\mathbb{R}^{l}$ its counterpart in the augmented state-reference space. Let $h:\mathcal{D}\to\mathbb{R}$ be continuously differentiable. We say $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ is a regular point of $h$ if $\partial h(\boldsymbol{\mathbf{x}})/\partial\boldsymbol{\mathbf{x}}\neq 0$. We say $c\in\mathbb{R}$ is a regular value of $h$ if every $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ such that $h(\boldsymbol{\mathbf{x}})=c$ is a regular point.
@@ 2.1 Control Invariance
Note that this property is sometimes called viability [28] or weak control invariance [27] because it doesn’t require existence and uniqueness of the trajectory. Nonetheless, it will simplify the presentation of the upcoming results. Assumptions for the existence and uniqueness of solutions will be presented separately. Central to the upcoming discussion is Nagumo’s theorem characterizing control invariance in terms of Bouligand’s tangent cone.
@@
A closed set $\mathcal{C}\subset\mathbb{R}^{n}$ is control invariant if and only if
@@
where $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is the tangent cone to $\mathcal{C}$ in $\boldsymbol{\mathbf{x}}$ [27, Definition 3.1].
While this characterization was groundbreaking, working with the tangent cone $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is cumbersome. Fortunately, when the set $\mathcal{C}$ is the 0-superlevel set of a continuously differentiable function, the tangent cone on the boundary is reduced to the tangent half-space shifted to the origin and we have the following characterization.
Theorem 2 ([30, Proposition 4.3.7])
Let $h:\mathcal{D}\to\mathbb{R}$ be a continuously differentiable function with regular value 0. The set $\mathcal{C}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}:h(\boldsymbol{\mathbf{x}})\geq 0\}$ is control invariant if and only if
@@
Note that [30, Proposition 4.3.7] is more general than Theorem 2: it allows the set $\mathcal{C}$ to be defined using multiple continuously differentiable functions. We use this fact in upcoming results.
@@ 2.2 Control Barrier Functions
Control barrier functions certify the control invariance of their superlevel set.
@@
A continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ is a control barrier function (CBF) if there exists a class $\mathscr{K}$ function $\alpha:[0,\infty)\to[0,\infty)$ such that
@@
Assume $\mathcal{U}\!\subset\!\mathbb{R}^{m}$ is compact. Let $\mathcal{C}\!\subset\!\mathbb{R}^{n}$, defined as in (5), be compact and let 0 be a regular value of $h$. The set $\mathcal{C}$ is control invariant if and only if $h$ is a CBF.
@@
is continuous on $\mathcal{D}$. Moreover, it can be shown that the set-valued map $C$ is both uppper- and lower-hemicontinuous. Thus, it follows by Berge’s maximum theorem [31], that the value function $\hat{\alpha}$ is continuous on $0,\infty)$. By construction, $\hat{\alpha}$ is non-decreasing. By Theorem 2, for any $\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$ such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq 0$. Thus, $\forall\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}=\mathcal{C}(0)$, $\xi(\boldsymbol{\mathbf{x}})\geq 0$. As a result,
@@
There always exists a class $\mathscr{K}$ function $\alpha:[0,\infty)\to[0,\infty)$ that upper-bounds $\hat{\alpha}$. Finally, let $\boldsymbol{\mathbf{x}}\in\mathcal{C}$ be given and note
@@
On the other hand, assume $h$ is a CBF, let $\alpha\in\mathscr{K}$ be such that (4) holds, and let $\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}$ be given. By compactness of $\mathcal{U}$, there exists $\boldsymbol{\mathbf{u}}\in\mathcal{U}$ such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq-\alpha\big(h(\boldsymbol{\mathbf{x}})\big)=0$, where the last equality holds because $h(\boldsymbol{\mathbf{x}})=0$. Since $\mathcal{C}$ is closed and 0 is a regular value of $h$, we can apply Theorem 2 to conclude $\mathcal{C}$ is control invariant.
@@
In this work we drop the compactness requirement on $\mathcal{U}$ because only the sufficiency statement in Lemma 1 is relevant and we will design CBFs for which (9) holds. Compactness of $\mathcal{C}$ is required for the necessity (control invariant $\mathcal{C}$ implies $h$ is a CBF) of Lemma 1, but not for the sufficiency ($h$ is a CBF implies $\mathcal{C}$ is control invariant) of Lemma 1.
@@
Consider $n_{c}\in\mathbb{N}$ continuously differentiable functions $h_{i}:\mathcal{D}_{i}\to\mathbb{R}$ and let $\mathcal{C}_{i}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}_{i}\mid h_{i}(\boldsymbol{\mathbf{x}})\geq 0\}\subset\mathbb{R}^{n}$, with $\mathcal{C}=\cap_{i=1}^{n_{c}}\mathcal{C}_{i}$. The functions $h_{i}$ are said to have the control-sharing property if there exists a class $\mathscr{K}$ function $\alpha$ satisfying $\forall\boldsymbol{\mathbf{x}}\in\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$, such that
@@
In this paper, we call a continuously differentiable function $h:\mathcal{D}\to\mathbb{R}^{n_{c}}$ whose channels $h_{i}$ have the control-sharing property a (vector-valued) control barrier function.
The functions $h_{i}$ in Definition 2.4 are deliberately not labeled as CBFs individually (as opposed to [9, Definition 2]) because such a requirement is unnecessarily limiting. In fact, if 0 is a regular value of each $h_{i}$, Lemma 1 also holds for the set $\mathcal{C}$ defined as the intersection of the $\mathcal{C}_{i}$ even if each $\mathcal{C}_{i}$ is individually not compact. The optimization-based policy $\beta$ can also be used with (vector-valued) CBFs with the following admissible input set
@@
Remark 2.5.
While we stated Definition 2.4 using a unique class $\mathscr{K}$ function $\alpha$ to simplify the presentation, it is equivalent to considering a different $\alpha_{i}\in\mathscr{K}$ for each $h_{i}$. Tuning each $\alpha_{i}$ allows the user to modulate the performance of the closed-loop system while retaining the safety guarantees as long as (13) is satisfied with $\alpha\to\alpha_{i}$.
As noted in the introduction, finding a CBF for arbitrary constraint sets $\mathcal{X}\subset\mathbb{R}^{n}$ is a challenging task. In fact, Lemma 1 states that, when $\mathcal{X},\mathcal{U}$ are compact, a CBF $h:\mathcal{D}\to\mathbb{R}$ satisfying $\mathcal{X}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}$ exists only if $\mathcal{X}$ is control invariant, which is generally not the case. In spite of this, many practitioners insist on imposing $\mathcal{C}=\mathcal{X}$ to construct a “candidate” CBF (i.e. a continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ that satisfies $\mathcal{X}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}$) and then tuning a class $\mathscr{K}$ function $\alpha$ on a case-by-case basis. Although this approach can yield good performance, it suffers from the possibility that $\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})=\emptyset$ for some $\boldsymbol{\mathbf{x}}\in\mathcal{X}$, which makes the controller $\beta$ ill-defined.